# Phase Retrieval by Quaternionic Reweighted Amplitude Flow on Image Reconstruction

### v1

@@ 3 Quaternionic Reweighted Amplitude Flow
The main goal of this section is to design the Reweighted Amplitude Flow algorithm for quaternion-valued signals (QRAF), which extends the Real/Complex Amplitude Flow (RAF) algorithm originally introduced in [8]. We organize the algorithm retaining the structure of the original RAF on purpose, however, it should keep in mind that the quaternion framework introduces non-commutativity.
We consider the quaternionic Gaussian measurement ensemble where the entries of ${\boldsymbol{A}}$ are i.i.d. drawn from
$\begin{split}\mathscr{N}_{\mathbb{H}}:=&\frac{1}{2}\left(\mathscr{N}(0,1)+\mathscr{N}(0,1)i\right.\\ &+\mathscr{N}(0,1)j+\left.\mathscr{N}(0,1)k\right),\end{split}$
denoted by ${\boldsymbol{A}}\sim\mathscr{N}_{\mathbb{H}}^{n\times d}$, yielding $\mathbb{E}(\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*})={\bf I}_{d}$.
@@ 3.1 Algorithm
$\min_{\boldsymbol{z}\in\mathbb{H}^{\,d}}F(\boldsymbol{z})=\frac{1}{n}\sum_{j=1}^{n}\left(|\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle|-\psi_{j}\right)^{2},$
where $\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle=\boldsymbol{\alpha}_{j}^{*}\,\boldsymbol{z}$ represents the quaternionic inner product, and $\psi_{j}=|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle|$ denotes the modulus of the quaternionic inner product between the known design vector $\boldsymbol{\alpha}_{j}$ and the unknown solution $\boldsymbol{x}$. We assume that $\|\boldsymbol{x}\|=1$.
QRAF begins with a reweighted initialization procedure, and subsequently refines the initial estimate $\boldsymbol{z}_{0}$ through a quaternion-based gradient descent.
@@ 3.1.1 Quaternionic Weighted Maximal Correlation Initialization
The importance of selecting an effective starting point is well-recognized for non-convex iterative algorithms in achieving the global optimum. The QRAF is not an exception.
The quaternionic weighted maximal correlation initialization consists of two steps. First, the norm of the true signal $\boldsymbol{x}$ is estimated easily as:
@@
$\begin{split}&\max_{\|\boldsymbol{z}\|=1}\frac{1}{|\mathcal{S}|}\sum_{j\in\mathcal{S}}\omega_{j}^{(0)}|\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle|^{2}\\ =\,&\boldsymbol{z}^{*}\left(\frac{1}{|\mathcal{S}|}\sum_{j\in\mathcal{S}}\omega_{j}^{(0)}\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{*}\right)\boldsymbol{z}.\end{split}$
@@ 3.1.2 Quaternionic Adaptively Reweighted Gradient Flow
where the adaptive weights are defined by
$\omega_{k}^{(t)}=\frac{1}{1+\beta/\left(|\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}|/|\boldsymbol{\alpha}_{k}^{*}\boldsymbol{x}|\right)},\qquad 1\leq k\leq n,$
in which the dependence on the iterate index $t$ is ignored for notational brevity. It should be keep in mind the multiplication order in the factor $\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}$ in (3.2) is crucial, otherwise the algorithm may fail.
The idea behind introducing this quaternionic gradient is as same as it in the conventional RAF in [8], which is to differentiate the contributions of various gradients to the overall search direction. A straightforward approach is thus to assign large weights to more reliable gradients and smaller weights to the less reliable ones. Therefore, the gradient is designed based on the ratio $\left|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{z}/\boldsymbol{\alpha}_{i}^{*}\boldsymbol{x}\right|$, which serves as a confidence score reflecting the reliability of the corresponding gradient and potentially indicating directions that lead to the true $\boldsymbol{x}$.
Taking a suitable step size $\eta$, the update rule is given by
$\boldsymbol{z}_{i+1}=\boldsymbol{z}_{i}-\eta\cdot\nabla\ell_{\rm rw}(\boldsymbol{z}_{i}).$
@@ 3.2 Linear Convergence Discussion
In this section, we present a partial analysis of the convergence of the QRAF algorithm. The primary challenge remains in establishing the local regularity condition, which is analogous to the difficulty encountered in the complex RAF algorithm, as discussed in the introduction. However, once this problem is resolved in the complex domain, we believe that extending the result to the quaternionic setting will not require significant additional effort.
@@
For the initial guess $\boldsymbol{z}_{0}$, produced by the spectrum method in Algorithm 1, it is found that the weighted initialization is effective, following a similar discussion in the complex case. Hence we omit the detailed proof.
@@
For an arbitrary $x\in\mathbb{H}^{d}$, consider the noiseless measurements $\psi_{i}=|\psi^{*}x|$, $1\leq i\leq n$. If $n\geq c_{0}|S|\geq c_{1}d$, then with probability exceeding $1-c_{3}e^{-c_{2}n}$, the initial guess $z_{0}$ obtained by the weighted maximal correlation method satisfies
${\rm dist}(z_{0},x)\leq\rho\|x\|$
for $\rho=1/10$. Here $c_{0},c_{1},c_{2},c_{3}>0$ are some absolute constants.
@@
$\begin{split}&{\rm Re}\left\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\right\rangle\\ \geq&\frac{\lambda}{2}{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\frac{\mu}{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}\end{split}$
holds for all $\boldsymbol{z}\in\mathbb{H}^{d}$ such that $\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|\leq\varepsilon\|\boldsymbol{x}\|$ for some constant $10<\varepsilon<$ 1. The ball given by $|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})|\leq\varepsilon\|\boldsymbol{x}\|$ is termed the basin of attraction in literature.
@@
For an arbitrary $\boldsymbol{x}\in\mathbb{H}^{d}$, consider $n$ noise-free measurements $\psi=|\boldsymbol{\alpha}_{j}^{*}\boldsymbol{x}|,1\leq j\leq n$. There exist some constants $c_{1},c_{2},c_{3}>0$, and $10<\nu<$ 1 such that the following holds with probability exceeding $1-c_{3}e^{-c_{2}n}$
@@
$\begin{split}&{\rm dist}^{2}(\boldsymbol{z}-\mu\cdot\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{x})\\ \leq\,&\|\boldsymbol{z}-\mu\nabla\ell_{{\rm rw}}(\boldsymbol{z})-\boldsymbol{x}\phi(\boldsymbol{z})\|^{2}\\ =\,&{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\mu^{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}\\ \,&-2\mu\cdot{\rm Re}\left\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\right\rangle\\ \leq\,&{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\mu^{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}\\ \,&-2\mu\left(\frac{\mu}{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}+\frac{\lambda}{2}{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})\right)\\ \leq\,&\left(1-\lambda\mu\right){\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x}),\end{split}$
@@
$\begin{split}\nabla\ell_{{\rm rw}}(\boldsymbol{z})&=\frac{1}{n}\sum_{k=1}^{n}\omega_{k}^{t}\left(1-\frac{y_{k}}{\left|{\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}}\right|}\right)\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*}z\\ &=\frac{1}{n}{\rm diag}(\boldsymbol{w})\boldsymbol{A}\boldsymbol{v},\end{split}$
@@
$\begin{split}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|&=\left\|\frac{1}{n}{\rm diag}(\boldsymbol{w})\boldsymbol{A}\boldsymbol{v}\right\|\\ &\leq\frac{1}{n}\|{\rm diag}(\boldsymbol{w})\|\cdot\|\boldsymbol{A}\|\cdot\|\boldsymbol{v}\|\\ &\leq\frac{1+\delta_{1}}{\sqrt{n}}\|\boldsymbol{v}\|,\end{split}$
where we have used the inequalities $\|{\rm diag}(\boldsymbol{w})\|\leq 1$ since $\omega_{i}\leq 1$ for all $1\leq i\leq n$ and $\|\boldsymbol{A}\|\leq(1+\delta_{1})\sqrt{n}$ for some constants $\delta_{1}>0$ (see e.g., [34]), assuming that $n/d$ is sufficiently large.
@@
$\begin{split}\|\boldsymbol{v}\|^{2}&\leq\sum_{k=1}^{n}(|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{z}|-|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{x}|)^{2}\\ &\leq\sum_{k=1}^{n}\|\boldsymbol{\alpha}_{i}^{*}(\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z}))\|^{2}\\ &\leq(1+\delta_{2})^{2}n\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|^{2},\end{split}$
@@
Combining these results, and taking $\delta>0$ larger than $(1+\delta_{1})(1+\delta_{2})-1$, the size of $\nabla\ell_{{\rm rw}}(\boldsymbol{z})$ can be bounded as
$\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|\leq(1+\delta)\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|,$
which holds with probability $1-e^{-c_{2}n}$, with a proviso that $n/d$ exceeds some numerical constant $c>0$.
@@
Finally, we still need to prove the following, which is open in the complex case as well. Once it is resolved in the complex setting, it will be possible to extend to the present quaternionic case without significant efforts.
@@
$\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{h}\rangle\geq c\|\boldsymbol{h}\|^{2}$
@@ 4.1 Incremental Algorithm: QIRAF–the critical one
In recent years, several incremental algorithms, such as Incremental Truncated Wirtinger Flow (ITWF) [27] and Incremental Reshaped Wirtinger Flow (IRWF) [16] have been developed. This further motivates the development of incremental or stochastic versions of QRAF, using mini-batches of measurements, referred to as incremental QRAF (QIRAF). The mini-batch QIRAF algorithm empolys the same initialization procedure as QRAF, and use a mini-batch of measurements for each gradient update. We describe it in Algorithm 2 below.

### v2

@@ 3 Quaternionic Reweighted Amplitude Flow
The main goal of this section is to design the Reweighted Amplitude Flow algorithm for quaternion-valued signals (QRAF), which extends the Real/Complex Amplitude Flow (RAF) algorithm originally introduced in [18]. We organize the algorithm retaining the structure of the original RAF on purpose, however, it should keep in mind that the quaternion framework introduces non-commutativity. We consider the quaternionic Gaussian measurement ensemble where the entries of ${\boldsymbol{A}}$ are i.i.d. drawn from $\mathscr{N}_{\mathbb{H}}=\frac{1}{2}\left(\mathscr{N}(0,1)+\mathscr{N}(0,1)i+\mathscr{N}(0,1)j+\mathscr{N}(0,1)k\right),$ denoted by ${\boldsymbol{A}}\sim\mathscr{N}_{\mathbb{H}}^{n\times d}$, yielding $\mathbb{E}(\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*})={\bf I}_{d}$.
@@ 3.1 Algorithm
$\min_{\boldsymbol{z}\in\mathbb{H}^{\,d}}F(\boldsymbol{z})=\frac{1}{n}\sum_{j=1}^{n}\left(|\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle|-\psi_{j}\right)^{2},$ (3.1)
where $\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle=\boldsymbol{\alpha}_{j}^{*}\,\boldsymbol{z}$ represents the quaternionic inner product, and $\psi_{j}=|\langle\boldsymbol{\alpha}_{j},\boldsymbol{x}\rangle|$ denotes the modulus of the quaternionic inner product between the known design vector $\boldsymbol{\alpha}_{j}$ and the unknown solution $\boldsymbol{x}$. We assume that $\|\boldsymbol{x}\|=1$. QRAF begins with a reweighted initialization procedure, and subsequently refines the initial estimate $\boldsymbol{z}_{0}$ through a quaternion-based gradient descent.
@@ 3.1.1 Quaternionic Weighted Maximal Correlation Initialization
The importance of selecting an effective starting point is well-recognized for non-convex iterative algorithms in achieving global optimum. QRAF is not an exception. The quaternionic weighted maximal correlation initialization consists of two steps. First, the norm of the true signal $\boldsymbol{x}$ is estimated easily as:
@@
$\begin{split}\max_{\|\boldsymbol{z}\|=1}\frac{1}{|\mathcal{S}|}\sum_{j\in\mathcal{S}}\omega_{j}^{(0)}|\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle|^{2}=\boldsymbol{z}^{*}\left(\frac{1}{|\mathcal{S}|}\sum_{j\in\mathcal{S}}\omega_{j}^{(0)}\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{*}\right)\boldsymbol{z}.\end{split}$
@@ 3.1.2 Quaternionic Adaptively Reweighted Gradient Flow
where the adaptive weights are defined by $\omega_{k}^{(t)}=1/\left(1+\beta/\left(|\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}|/|\boldsymbol{\alpha}_{k}^{*}\boldsymbol{x}|\right)\right)$, for $1\leq k\leq n,$ in which the dependence on the iterate index $t$ is ignored for notational brevity. It should be keep in mind the multiplication order in the factor $\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}$ in (3.3) is crucial, otherwise the algorithm may fail.
The idea behind introducing this quaternionic gradient is as same as it in the conventional RAF in [18], which is to differentiate the contributions of various gradients to the overall search direction. A straightforward approach is thus to assign large weights to more reliable gradients and smaller weights to the less reliable ones. Therefore, the gradient is designed based on the ratio $\left|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{z}/\boldsymbol{\alpha}_{i}^{*}\boldsymbol{x}\right|$, which serves as a confidence score reflecting the reliability of the corresponding gradient and potentially indicating directions that lead to the true $\boldsymbol{x}$. Taking a suitable step size $\eta$, the update rule is given by $\boldsymbol{z}_{i+1}=\boldsymbol{z}_{i}-\eta\cdot\nabla\ell_{\rm rw}(\boldsymbol{z}_{i})$.
@@ 3.2 Linear Convergence Discussion
In this section, we present a partial analysis of the convergence of the QRAF algorithm. The primary challenge remains in establishing the local regularity condition, which is analogous to the difficulty encountered in the complex RAF algorithm. However, once this problem is resolved in the complex domain, we believe that extending the result to the quaternionic setting will not require significant additional effort.
@@
For the initial guess $\boldsymbol{z}_{0}$ by spectrum method in Algorithm 1, it is found that the weighted initialization is effective, following a similar discussion in the complex case. Hence we omit the detailed proof.
@@
For an arbitrary $x\in\mathbb{H}^{d}$, consider the noiseless measurements $\psi_{i}=|\psi^{*}x|$, $1\leq i\leq n$. If $n\geq c_{0}|S|\geq c_{1}d$, then with probability exceeding $1-c_{3}e^{-c_{2}n}$, the initial guess $z_{0}$ obtained by the weighted maximal correlation method satisfies ${\rm dist}(z_{0},x)\leq\rho\|x\|$ for $\rho=1/10$. Here $c_{0},c_{1},c_{2},c_{3}>0$ are some absolute constants.
@@
$\begin{split}&{\rm Re}\left\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\right\rangle\geq\frac{\lambda}{2}{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\frac{\mu}{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}\end{split}$
holds for all $\boldsymbol{z}\in\mathbb{H}^{d}$ such that $\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|\leq\varepsilon\|\boldsymbol{x}\|$ for some constant $10<\varepsilon<$ 1. The ball given by $\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|\leq\varepsilon\|\boldsymbol{x}\|$ is termed the basin of attraction in literature.
@@
For an arbitrary $\boldsymbol{x}\in\mathbb{H}^{d}$, consider $n$ noise-free measurements $\psi=|\boldsymbol{\alpha}_{j}^{*}\boldsymbol{x}|,1\leq j\leq n$. There exist some constants $c_{1},c_{2},c_{3}>0$, and $10<\nu<$ 1 such that the following holds with probability exceeding $1-c_{3}e^{-c_{2}n}$,
@@
$\begin{split}{\rm dist}^{2}(\boldsymbol{z}-\mu\cdot\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{x})\leq\,&\|\boldsymbol{z}-\mu\nabla\ell_{{\rm rw}}(\boldsymbol{z})-\boldsymbol{x}\phi(\boldsymbol{z})\|^{2}\\ =\,&{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\mu^{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}-2\mu\cdot{\rm Re}\left\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\right\rangle\\ \leq\,&{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})+\mu^{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}-2\mu\left(\frac{\mu}{2}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|^{2}+\frac{\lambda}{2}{\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x})\right)\\ \leq\,&\left(1-\lambda\mu\right){\rm dist}^{2}(\boldsymbol{z},\boldsymbol{x}),\end{split}$
@@
$\begin{split}\nabla\ell_{{\rm rw}}(\boldsymbol{z})&=\frac{1}{n}\sum_{k=1}^{n}\omega_{k}^{t}\left(1-\frac{y_{k}}{\left|{\boldsymbol{\alpha}_{k}^{*}\boldsymbol{z}}\right|}\right)\boldsymbol{\alpha}_{k}\boldsymbol{\alpha}_{k}^{*}z=\frac{1}{n}{\rm diag}(\boldsymbol{w})\boldsymbol{A}\boldsymbol{v},\end{split}$
@@
$\begin{split}\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|&=\left\|\frac{1}{n}{\rm diag}(\boldsymbol{w})\boldsymbol{A}\boldsymbol{v}\right\|\leq\frac{1}{n}\|{\rm diag}(\boldsymbol{w})\|\cdot\|\boldsymbol{A}\|\cdot\|\boldsymbol{v}\|\leq\frac{1+\delta_{1}}{\sqrt{n}}\|\boldsymbol{v}\|,\end{split}$
where we have used the inequalities $\|{\rm diag}(\boldsymbol{w})\|\leq 1$ since $\omega_{i}\leq 1$ for all $1\leq i\leq n$ and $\|\boldsymbol{A}\|\leq(1+\delta_{1})\sqrt{n}$ for some constants $\delta_{1}>0$, assuming that $n/d$ is sufficiently large.
@@
$\begin{split}\|\boldsymbol{v}\|^{2}&\leq\sum_{k=1}^{n}(|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{z}|-|\boldsymbol{\alpha}_{i}^{*}\boldsymbol{x}|)^{2}\leq\sum_{k=1}^{n}\|\boldsymbol{\alpha}_{i}^{*}(\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z}))\|^{2}\leq(1+\delta_{2})^{2}n\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|^{2},\end{split}$
@@
Combining these results, and taking $\delta>0$ larger than $(1+\delta_{1})(1+\delta_{2})-1$, the size of $\nabla\ell_{{\rm rw}}(\boldsymbol{z})$ can be bounded as $\|\nabla\ell_{{\rm rw}}(\boldsymbol{z})\|\leq(1+\delta)\|\boldsymbol{z}-\boldsymbol{x}\phi(\boldsymbol{z})\|,$ which holds with probability $1-e^{-c_{2}n}$, with a proviso that $n/d$ exceeds some numerical constant $c>0$.
@@
Finally, we still need to prove the following, which is open in the complex case as well.
@@
${\rm Re}\langle\nabla\ell_{{\rm rw}}(\boldsymbol{z}),\boldsymbol{h}\rangle\geq c\|\boldsymbol{h}\|^{2}$ (3.8)
@@
Such a bound, established in the real number setting, relies heavily on the fact for $j=1,\cdots,n$, the value of $\text{sign}\langle\boldsymbol{\alpha}_{j},\boldsymbol{z}\rangle$ is always either $-1$ or $1$, which can only hold in the real number field. A straightforward generalization of proof to the complex domain is very challenging, as realized by several researchers, see e.g.,[22] and [23]. However, once resolved in complex setting, the problem can be further extended to the present quaternionic case using existing techniques in quaternionic analysis [20]. Alternatively, to circumvent this difficulty, perturbed amplitude flow model generalizing [22] is introduced in Section 5. Within this framework, the desired lower bound (3.8) can be established rigorously. Finally, we believe that the global convergence is ensured by a similar (but more subtle) favorable benign geometric landscape, as established in [24].
@@ 4.1 Incremental Algorithm: QIRAF–the critical one
Recently, incremental algorithms, such as Incremental Reshaped Wirtinger Flow (IRWF, see e.g. [25]) , have been developed. This further motivates the development of incremental or stochastic versions of QRAF, using mini-batches of measurements, referred to as incremental QRAF (QIRAF). The mini-batch QIRAF algorithm empolys the same initialization procedure as QRAF, and use a mini-batch of measurements for each gradient update. We describe it in Algorithm 2 below.